OpenStudy (anonymous):

How to Factor (x^6)-(y^6) using difference of squares?

OpenStudy (anonymous):

\[x^6-y^6\]

OpenStudy (aum):

\[ x^6 - y^6 = (x^3)^2 - (y^3)^2 = ? \]

OpenStudy (aum):

\[ a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

I got \[(x^3-y^3)(x^3+y^3) \] but how would factor this even further?

OpenStudy (anonymous):

is that wrong?

OpenStudy (aum):

That is correct. That is the first step. Then you can factor further by using the following identities: \[ a^3 - b^3 = (a-b)(a^2 + ab + b^2) \\ a^3 + b^3 = (a+b)(a^2 - ab + b^2) \]

OpenStudy (anonymous):

oh okay the difference of cubes, thats whats i was missing i think,

OpenStudy (anonymous):

wait hold on, how would i go about plugging \[(x^3-y^3)(x^3+y^3)\]

OpenStudy (anonymous):

into the identity

OpenStudy (freckles):

replace a with x and replace y with b

OpenStudy (anonymous):

so I just do this: \[(x-y)^3(x+y)^3=((x-y)-(x+y)((x-y)^2+(x-y)(x+y)+(x+y))\]

OpenStudy (anonymous):

and then simplify

OpenStudy (freckles):

didn't you have \[(x^3-y^3)(x^3+y^3)?\]

OpenStudy (anonymous):

yes

OpenStudy (freckles):

use what aum gave you to factor more \[a^3 - b^3 = (a-b)(a^2 + ab + b^2) \\ a^3 + b^3 = (a+b)(a^2 - ab + b^2)\]

OpenStudy (freckles):

like you can factor x^3-y^3 more by saying (x-y)(x^2+xy+y^2) now you need to factor x^3+y^3

OpenStudy (anonymous):

Thanks guys I got it :D